# Inferring additional physics through unmodelled signal reconstructions

### v2

@@ III.2.1 Reconstructed Waveforms
We perform the cWB-waveform reconstructions for all injections mentioned in the PE section. The details about cWB reconstruction are discussed in Section II.2. These analyses were performed in Gaussian noise with Advanced LIGO and Virgo sensitivity.
Fig 3 shows the reconstructed whitened strain for parameter estimation of select hybrids. The orange curve represents the reconstructed whitened waveform for a hybrid injection by cWB for the LIGO Livingston detector. The light gray shaded regions denote the 90% credible intervals derived from the waveform reconstruction of the PE samples, and the black curve corresponds to the median of these reconstructed waveforms. The top, middle and bottom rows represent results from selected hybrids with , respectively. The left and right columns refer to zero and maximum eccentricity hybrid injection, respectively, for a given mass ratio. It is well understood that eccentricity influences the signal duration; for a fixed-mass binary, the signal duration decreases as eccentricity increases. This trend is observable in Fig 3, moving from left to right.
We observe that for hybrids with no eccentricity (left column of Fig. 3), the reconstructed signal lies within the 90% credible interval (shaded gray region). However, for eccentric hybrid injections (right column of Fig. 3), discrepancies arise between the 90% reconstruction band and the injected reconstruction. These discrepancies become more pronounced as eccentricity increases and are observed in both the time and frequency domains. For hybrids with higher eccentricity (e.g., BBH_1361, BBH_1363), the reconstructed strain not only falls significantly outside the gray region but also shows a considerable phase deviation from the median of the PE samples.
@@ III.2.2 Figures of Merit: Overlap Distribution
In this section, we present the results of the figure-of-merit analysis conducted for the injected signals. Overlap factors were computed using Equation (1) for all simulated signals, following the methodology outlined in the flow chart. Fig. 4 illustrates the overlap distributions for on-source (orange) and null (blue) cases across various cWB reconstruction analyses for . The vertical dashed lines denote the median for on-source and null distributions respectively.
@@
We observe that the null histogram exhibits a relatively consistent distribution of overlap values across different cases. However, this behaviour changes for high eccentricity () case, most likely due to a loss in low-energy pixels during the reconstruction process. Significant deviations in the overlap between the two histograms can be identified with increasing eccentricity. Specifically, the overlap refers to the common region enclosed by the two histograms within their 90% area about the median. Beyond a certain eccentricity threshold, the overlap diminishes to nearly zero, indicating a substantial difference between the PE sample cWB waveform reconstruction and the hybrid injection reconstruction.
To quantify this behaviour, we introduce a new metric that measures the difference between the medians of the two histograms (Eq. 2) as a function of eccentricity. This metric provides a clearer indication of how waveform reconstruction consistency degrades with increasing eccentricity. We present a figure of merit Fig. 5 for different hybrids. Each data point in the scatter plot corresponds to a specific eccentricity and mass ratio, capturing trends and variations across these parameters. As eccentricity increases, the value of increases, indicating that the eccentricity feature is observable in . This trend remains consistent across different mass ratios. We have plotted the figure of merit with error bars where, errors in the median shift are calculated using standard error propagation method. For q = 1, 2 and 3, the maximum errors in the median shift corresponding to the highest eccentricity are 0.024, 0.018 and 0.020 respectively. We have also calculated the relative shifts in for eccentric injections with respect to the circular one. The values for the relative shifts are 15.9, 8.45 and 137.17 corresponding to highest eccentricity for q = 1, 2 and 3 respectively. These deviations represent the combined contributions of biases in CBC parameter estimation and effects from cWB waveform reconstruction. It is important to note that this analysis was conducted in Gaussian noise; for real noise scenarios, these results will be influenced by noise fluctuations.
@@ IV Conclusion and future directions
Our results can be divided into two parts: parameter estimation analysis and waveform reconstruction study (see Fig. 1 for a flow chart describing the methodology). We inject hybrid signals of total mass 40 into Gaussian noise and perform PE using a quasi-circular waveform model, IMRPhenomXAS [120]. The violin plot in Fig. 2 shows the effect of ignoring eccentricity greater than 0.1 at 20Hz () for (1,2,3) as the recovered chirp mass is outside the 90% credible interval. Similar biases are also noticed in recovering other intrinsic parameters such as total mass and mass ratio with increasing eccentricity. Additionally, in the waveform reconstruction analysis (see Fig. 3) we observe that for the reconstructed waveform is inconsistent with the 90% credible interval and is out of phase. We compute the overlap factor for on-source and null distributions discussed in section II.3. For the on-source analysis, we find that with increasing eccentricity, the overlap factor reduces sharply and reaches 75% for . As a result, the on-source distribution deviates from the null distribution, which is also evident from Fig 5.
Thus, we can conclude that the orbital eccentricity of a binary system, which is a sub-dominant effect, can be isolated even in cases where eccentric waveforms are not readily accessible. The methodology is so robust that it can be employed to infer the presence of any sub-dominant physical effect, such as the presence of precession effects in the signal due to spin and/or higher-order modes apart from eccentricity. In future work, we plan to extend the methods presented here to analyse real events, which, in principle, can include a combination of these sub-dominant effects.
@@ Acknowledgements.
We thank Juan Calderón Bustillo for his useful comments on the manuscript. RD thanks the Gravitation and Cosmology Group at IIT Madras for numerous insightful discussions. RD also thanks Tanmaya Mishra for helping fix a few issues with the cWB pipeline installation. DJ acknowledges the Science and Technology Facilities Council (STFC) for support through grants ST/V005618/1 and ST/Y004272/1. G.V. acknowledges the support of NSF under grant PHY-2207728. S.K. acknowledges the support of NSF under grants PHY 2110060 and PHY-2409372. I.B acknowledges the support of NSF under grants PHY-2309024. C.K.M. acknowledges the support of SERB’s Core Research Grant No. CRG/2022/007959. Computations were performed on CIT, LHO, and LLO clusters provided by the LIGO Laboratory and supported by the National Science Foundation Grants PHY-0757058 and PHY-0823459. This material is based upon work supported by NSF’s LIGO Laboratory, which is a major facility fully funded by the National Science Foundation. We used the following software packages: LALSuite [122], bilby [125], bilby_pipe [126], PyCBC [129], NumPy [130], PESummary [131], Matplotlib [132], Seaborn [133], jupyter [134], and dynesty [124]. This document has LIGO preprint No. LIGO-P2400590.Appendix A Priors used for parameter estimation
The sample space for parameter estimation includes the following parameters: inverse mass ratio (),333We use the word inverse here to indicate that this is inverse of the mass ratio we have used throughout the paper. Please note that bilby uses the term mass_ratio for this. chirp mass (), luminosity distance (), inclination angle (), geocentric time (), phase angle (), right ascension (), declination (), and polarization angle (). We have put constraint on component masses as [5,50]. The priors for all the parameters are given in Table 1.

### v3

@@ III.2.1 Reconstructed waveforms
We perform the cWB-waveform reconstructions for all injections mentioned in the PE section. The details about cWB reconstruction are discussed in Sec. II.2. These analyses were performed in Gaussian noise with Advanced LIGO and Virgo sensitivity.
Figure 3 shows the reconstructed whitened strain for parameter estimation of selected hybrids. The orange curve represents the reconstructed whitened waveform for a hybrid injection by cWB for the LIGO Livingston detector which also acts as a reference detector for our analysis. The light gray shaded regions denote the 90% credible intervals derived from the waveform reconstruction of the PE samples, and the black curve corresponds to the median of these reconstructed waveforms. The top, middle and bottom rows represent results from selected hybrids with , respectively. The left and right columns refer to zero and maximum eccentricity hybrid injection, respectively, for a given mass ratio. It is well understood that eccentricity influences the signal duration; for a fixed-mass binary, the signal duration decreases as eccentricity increases. This trend is observable in Fig 3, moving from left to right.
We observe that for hybrids with no eccentricity (left column of Fig. 3), the reconstructed signal lies within the 90% credible interval (shaded gray region). However, for eccentric hybrid injections (right column of Fig. 3), discrepancies arise between the 90% reconstruction band and the injected reconstruction. These discrepancies become more pronounced as eccentricity increases and are observed in both the time and frequency domains. For hybrids with higher eccentricity (e.g., HYB:SXS:BBH:1361, HYB:SXS:BBH:1363), the reconstructed strain not only falls significantly outside the gray region but also shows a considerable phase deviation from the median of the PE samples.
@@ III.2.2 Figures of merit: Overlap distribution
In this section, we present the results of the figure-of-merit analysis conducted for the injected signals. Overlap factors were computed using Eq. (1) for all simulated signals, following the methodology outlined in the flow chart. Figure. 4 illustrates the overlap distributions for on-source (orange) and null (blue) cases across various cWB reconstruction analyses for . The vertical dashed lines denote the median for on-source and null distributions respectively.
@@
We observe that the null histogram exhibits a relatively consistent distribution of overlap values across different cases. However, this behaviour changes for high eccentricity () case, most likely due to a loss in low-energy pixels during the reconstruction process. Significant deviations in the shared region between the two histograms can be identified with increasing eccentricity. Beyond a certain eccentricity threshold, the area of the shared region diminishes to nearly zero, indicating a substantial difference between the PE sample cWB waveform reconstruction and the hybrid injection reconstruction.
To quantify this behaviour, we introduce a new metric that measures the difference between the medians of the two histograms (Eq. 3) as a function of eccentricity. This metric provides a clearer indication of how waveform reconstruction consistency degrades with increasing eccentricity. We present a figure of merit Fig. 5 for different hybrids. Each data point in the scatter plot corresponds to a specific eccentricity and mass ratio, capturing trends and variations across these parameters. As eccentricity increases, for the value of increases from 0.01 () to 0.2 () indicating that the eccentricity feature is observable in (Fig 4). For the value of increases from 0.02 () to 0.14 () and from 0.001 () to 0.17 (), respectively. As can be verified from Figure 5, a criterion of 4% for upper bounds on values (shown as black-dashed line) eliminates cases with eccentricity values (estimated at 20 Hz for a binary of total mass of using the gw-eccentricity package of Ref [122]) smaller than 0.17 as potential eccentric sources. As indicated in Figure 1, we use this to classify the source as a potential eccentric source or otherwise for follow up studies. We have plotted the figure of merit with error bars where, errors in the median shift are calculated using standard error propagation method (Eq. 4). These deviations represent the combined contributions of biases in CBC parameter estimation and effects from cWB waveform reconstruction. It is important to note that this analysis was conducted in Gaussian noise; for real noise scenarios, these results will be influenced by noise fluctuations.
@@ IV Conclusion and future directions
Our results can be divided into two parts: parameter estimation analysis and waveform reconstruction study (see Fig. 1 for a flowchart describing the methodology). We inject hybrid signals of total mass 40 into Gaussian noise and perform PE using a quasicircular waveform model, IMRPhenomXAS [123]. The violin plot in Fig. 2 shows the effect of ignoring eccentricity greater than 0.17 at 20Hz () for = 1,2 and 3 as the recovered chirp mass is outside the 90% credible interval. Similar biases are also noticed in recovering other intrinsic parameters such as total mass and mass ratio with increasing eccentricity. Additionally, in the waveform reconstruction analysis (see Fig. 3) we observe that for the reconstructed waveform is inconsistent with the 90% credible interval and is out of phase. We compute the overlap factor for on-source and null distributions discussed in Sec II.3. For the on-source analysis, we find that with increasing eccentricity, the overlap factor reduces sharply and reaches 75% for . As a result, the on-source distribution deviates from the null distribution, which is also evident from Fig 5.
Thus, we can conclude that the orbital eccentricity of a binary system, which is a subdominant effect, can be isolated even in cases where eccentric waveforms are not readily accessible. The methodology is so robust that it can be employed to infer the presence of any sub-dominant physical effect, such as the presence of precession effects in the signal due to spin and/or higher-order modes apart from eccentricity. In future work, we plan to extend the methods presented here to analyse real events, which, in principle, can include a combination of these subdominant effects.
@@ Acknowledgements.
We thank Juan Calderón Bustillo for his useful comments on the manuscript. RD thanks the Gravitation and Cosmology Group at IIT Madras for numerous insightful discussions. R.D. also thanks Tanmaya Mishra for helping fix a few issues with the cWB pipeline installation. D.J. acknowledges the Science and Technology Facilities Council (STFC) for support through grants ST/V005618/1 and ST/Y004272/1. G.V. acknowledges the support of NSF under grant PHY-2207728. S.K. acknowledges the support of NSF under grants PHY 2110060 and PHY-2409372. I.B acknowledges the support of NSF under grants PHY-2309024. C.K.M. acknowledges the support of SERB’s Core Research Grant No. CRG/2022/007959. Computations were performed on CIT, LHO, and LLO clusters provided by the LIGO Laboratory and supported by the National Science Foundation Grants PHY-0757058 and PHY-0823459. This material is based upon work supported by NSF’s LIGO Laboratory, which is a major facility fully funded by the National Science Foundation. We used the following software packages: LALSuite [125], bilby [128], bilby_pipe [129], PyCBC [132], NumPy [133], PESummary [134], Matplotlib [135], Seaborn [136], jupyter [137], and dynesty [127]. This document has LIGO preprint No. LIGO-P2400590.Appendix A PRIORS USED FOR PARAMETER ESTIMATION
The sample space for parameter estimation includes the following parameters: inverse mass ratio (),555We use the word inverse here to indicate that this is inverse of the mass ratio we have used throughout the paper. Please note that bilby uses the term mass_ratio for this. chirp mass (), luminosity distance (), inclination angle (), geocentric time (), phase angle (), right ascension (), declination (), and polarization angle (). We have put constraint on component masses as [5,50]. The priors for all the parameters are given in Table 1.